# FPGA-Enabled Machine Learning Applications in Earth Observation: A Systematic Review

### v1

@@ A Large Catalog of Computing Platforms
While many hardware platforms are available to accelerate AI inference on the edge, only a few COTS systems support heavier models like DNNs. Examples include NVIDIA’s Jetson Nano GPU (NVIDIA, 2019), Google’s Coral TPU (Google, 2019), and Intel’s Myriad 2 VPU (Intel, 2016), all presenting solid alternatives for terrestrial edge computing. However, their susceptibility to radiation limits their use in space applications, where FPGAs offer a more resilient alternative (Rapuano et al., 2021). While floating-point operations on FPGAs may exhibit comparable—or worse, due to lower clock frequency—performance to other embedded platforms, the true advantage of FPGAs emerges when leveraging fixed-point, typically 8-bit, or even binary (1-bit) operations (Wang et al., 2020). This capability, often not fully supported or efficiently implemented on other hardware platforms, allows FPGAs to achieve significant performance and efficiency gains.
Through the motivations of the surveyed studies, and in line with the conclusions of Wang et al. (2020) and Lentaris et al. (2018), a generic guideline for hardware platform selection, considering development effort and performance targets, can be formulated. GPUs, TPUs, and VPUs offer quicker paths to achieving initial performance, especially for computationally complex DNNs (Bayer et al., 2023). FPGAs, however, promise superior performance and efficiency if more development effort and hardware expertise are invested in custom accelerator design and optimization. For applications demanding ultimate performance and willing to bear the highest development costs and longest design cycles, custom ASICs represent the pinnacle, albeit with reduced flexibility (Amara et al., 2006). In the following sections, we look at the technical aspects of the surveyed studies, exploring the methodologies and techniques employed to implement and optimize ML models on FPGAs for RS applications.
AMD FPGAs overwhelmingly dominate the surveyed hardware, comprising nearly all devices ($98\%$). The preferred FPGAs are SoC boards, particularly the Zynq family ($57\%$). Most boards are equipped to run DNNs ($61\%$ with $>500$ DSPs), and almost all devices have accelerated at least one DL model in an experiment. Reconfigurability, power efficiency, and radiation concerns are the primary motivations for choosing FPGAs over other DNN-friendly COTS accelerators. 2.1 TAKEAWAYS
@@ 4. From AI Models to FPGA Deployment: Design Strategies and Challenges
Before deploying a model on an FPGA, it typically requires adaptation and compression. These optimization steps are closely linked to hardware design decisions—discussed in  2.1 and  2.1. In  2.1, we introduce key network compression techniques and discuss their impact and requirements. Since traditional ML methods are already lightweight, we primarily discuss optimizations for computationally intensive models, like NNs.
@@
Most NNs include a feature extractor, or backbone, responsible for refining and compressing the input’s essential information. Extracting rich, high-quality features often relies on large and deep backbones, which come with high computational costs. To lighten these costs, designers can exploit lightweight backbones. For instance, Yang et al. (2023) replace YOLOv3’s DarkNet53 backbone with GhostNet (Han et al., 2020), reducing the number of learnable parameters by a factor of $3$. Similarly, Li et al. (2023) swap YOLOv4’s CSPDarknet53 backbone for MobileNeXt (Zhou et al., 2020), lowering the parameter count from $60.08$M to $36.44$M, with only a $1.65\%$ mAP drop—compensated via other customizations.
As discussed in  2.1, the existence of many architectures make selection a challenging task. Sabogal and George (2021) compare four DL segmentation models, showing that ENet (Paszke et al., 2016) outperforms U-Net while being 23.81×23.81$\times$23.81 × more computationally efficient and requiring 20.56×20.56$\times$20.56 × less memory. Similarly, Yang et al. (2022) compare three popular lightweight backbones: MobileNetv1, MobileNetv2, and SqueezeNet (Iandola et al., 2016), concluding that MobileNetv1 achieves the best trade-off for their SAR ship detection application. MobileNetv1 (Howard et al., 2017) also introduces a width multiplier for layer-wise scaling, leveraged by Suh et al. (2021) to explore different deployment scenarios in terms of accuracy and FPGA resource use. Beyond swapping architectures, Kim et al. (2024) propose a multi-phase cloud coverage detection pipeline where images first undergo a uniformity check and classification via a small CNN (TriCloudnet) before full segmentation with U-Net. Such workflow avoids the computation of cloud-clear or overcast images, reducing processing time and power consumption by $48.7\%$. When sacrificing some feature representation capabilities is manageable, lightweight and multi-phase approaches highlight that questioning a model’s depth and width can result in significant computational savings.
@@
[…]
@@
[…]
@@
One of FPGA’s key computational advantages over alternative hardware platforms, like GPUs, lies in custom datatype operations. Quantization reduces numerical precision by converting full-precision 32-bit floating-point (fp32) parameters into lower-precision formats such as 16-bit fixed-point (i16), 8-bit integer (i8), or even binary (b). Low-bandwidth datatypes enable highly efficient arithmetic operations, particularly on FPGA logic, where binary multiplications can be performed using simple XOR operations. While quantization reduces computational cost and memory footprint, it inevitably perturbs model parameters, [^13] Several factors influence the effectiveness of a quantization scheme, such as its uniformity, symmetry, and granularity. For a more detailed discussion on quantization techniques, we recommend the comprehensive survey of Gholami et al. (2021). shifting them from their convergence point and causing accuracy degradation compared to full-precision implementations (Gholami et al., 2021). For instance, Sabogal and George (2021) report an average accuracy drop of $1\%$ across four segmentation models. Similarly, in object detection, Zhao et al. (2023) observe a $0.42\%$ drop in mAP, while Ni et al. (2023) report a classification accuracy reduction of $0.03\%$ for VGG16 and $0.06\%$ for ResNet-34. Although the quantization-induced loss rarely exceeds $1\%$, it can be unacceptable in scenarios requiring high model reliability.
To mitigate accuracy degradation, designers may opt to re-train the network, effectively choosing one of two standard approaches: Post-Training Quantization (PTQ) or Quantization-Aware Training (QAT). PTQ, sometimes called direct quantization, is straightforward to implement as it applies quantization directly after training. However, PTQ’s simplicity can cause significant accuracy degradation, especially in models with a wide parameter distribution (Jacob et al., 2018). For instance, Wei et al. (2019) observe over $50\%$ accuracy loss across all tested bit widths when applying PTQ to a LeNet-5 for MSTAR classification. In contrast, QAT simulates low-precision arithmetic already during training, resulting in better robustness to quantization noise (Pitonak et al., 2022), albeit at a higher computational cost (Yu et al., 2024).
Beyond PTQ and QAT, mixed-precision quantization assigns different bit widths to different sections of the model. This method preserves accuracy by keeping sensitive layers at higher precision while applying aggressive quantization to less critical layers. For example, Yang et al. (2022) identify inefficient layers by estimating layer-wise latency and resource consumption. Rapuano et al. (2021) take this idea further by manually assigning custom bit widths to each input, filter, and output, achieving a $48\%$ memory footprint reduction for CloudScout. Alternatively, some approaches like Zhang et al. (2020) mitigate accuracy loss by selectively retaining fp32 precision for activations while quantizing only the weights. Similarly, Wei et al. (2019) introduce a symmetric QAT method that quantizes weights but preserves normalization and activation in fp32 format. The resulting computation overhead is motivated by the need to preserve the model’s accuracy.
Quantization is indispensable to achieve FPGA designs that are competitive with higher-frequency computing platforms. While PTQ is often sufficient to achieve 8-bit quantization with minimal accuracy degradation (Nagel et al., 2021), QAT is preferable when maintaining accuracy is critical. Mixed-precision quantization further optimizes efficiency by balancing accuracy and computational cost.
@@
All the methods discussed above effectively optimize NNs, but their combined use often yields the best results. For example, Zhao et al. (2023) analyze in an ablative fashion the parameters and accuracy reduction of each subsequent technique applied. The resulting figure allows to clearly track which method provides the highest parameter reduction—backbone replacement in their case—and which method helps recover lost accuracy—here, data augmentation. Similarly, Yang et al. (2023) follow a structured design flow to progressively refine YOLOv4. They first replace its backbone with GhostNet (Han et al., 2020) (leveraging depth-wise separable convolutions), then apply channel pruning to drastically reduce parameters. To restore accuracy, they incorporate Knowledge Distillation (KD) (Hinton et al., 2015), before finally quantizing the model. This multi-step process results in a model with just $\approx 10\%$ of the original parameters and $5\%$ of the memory footprint. A similar approach is presented by Han et al. (2016), who combine pruning, quantization, and Huffman coding (weight clustering) to achieve a 35×35$\times$35 × storage reduction on AlexNet and 49×49$\times$49 × on VGG16, all without accuracy loss. Ultimately, the effectiveness of these optimization techniques depends not only on model compression but also on how they align with FPGA accelerator design, which we explore next.
To reduce computational load, $26\%$ of DNNs leverage lightweight backbones or efficient architectures, primarily MobileNets ($43\%$). Out of all DL experiments, $91\%$ use quantization, mostly to int8888 format, and $26\%$ leverage pruning to compress their model. Fine granularity mixed precision quantization presents significant opportunities for further network compression. 2.1 TAKEAWAYS
@@
Implementing NNs on CPUs and GPUs typically involves using stable and mature software frameworks and drivers. In contrast, automatic deployment frameworks for FPGAs have not seen the same level of adoption—most implementations depend on manual frameworks and varying design patterns. To support future work in selecting suitable implementation frameworks and design patterns, we provide a dedicated second taxonomy in Table 3. It presents key metrics relevant to the respective FPGA implementation, organized into three column groups. First, Implementation Choices includes the implementation framework ’Impl.’, the model family ’Fam.’, the design pattern ’P’ (S for Specific, F for Flexible), the ’FPGA’ product, and the ’Model Name’. Second, Design Metrics covers the precision post quantitization ’Prec.’, the model complexity ’C[OP], and the memory footprint [MB] (: model parameters are located on-chip, and not loaded from external memory). Additionally, we report the DSP ’D[%] and BRAM ’B[%] resource utilization. Finally, Performance Metrics attempts to compare different designs with each other. In particular, we report the operating frequency in [MHz], computational throughput [^14] See Section 6.3 for a discussion about the confusion between [GOP], [GOPs], and [GOP/s]. ’T[GOP/s], power consumption ’P[W], computational efficiency ’T/P’, Latency ’L[s], and temporal throughput in [FPS].
@@
Manual HDL CNN S (Ieracitano et al., 2024) 7Z045 HO-ShipNet i16 - - 44 44 - - 1.9 - 1.0e-3 -
(Rapuano et al., 2021) ZU7EV CloudScout i? - 13.3 67 46 115 - 3.4 - 1.4e-1 -
(Vitolo et al., 2024) 7A35T CNN f32* - - - - 36 1.1 0.1 9.8 - -
(Yang et al., 2023) 7Z020 Ghost-YOLOS i16 8.41G 12.6 58 60 150 29.5 3.0 9.9 3.2e-1 -
(Zhang et al., 2020) 7VX690T Q-IORN i8 - 121.5 21 27 200 209.6 6.3 33.2 1.5e-1 6.8
(Zhang et al., 2023b) 7VX690T A2NN i8,f32 14.9G - 9 86 200 3047 8.3 368.4 4.9e-3 203.2
F (He et al., 2023) ZU15EG 2D CNN i8 684 K 1.2 5 71 100 7.1 8.4 0.8 9.7e-5* -
ZU15EG 3D CNN i8 4.23M 0.1 9 76 100 3.8 8.4 0.5 1.1e-3* -
ZU15EG HybridSN i8 101 M 20.5 37 80 100 13.2 8.4 1.6 7.7e-3* -
(Li et al., 2019) 7Z100 CBFF-SSD i16 - - 57 60 200 452.8 19.5 23.2 4.3e-2 -
(Li et al., 2020a) ZU3EG AP2D-Net i(mix) - - 80 75 300 130.2 5.6 23.3 3.3e-2 30.5
(Ni et al., 2023) 7VX690T Impr. YOLOv2 i8,i32 379 G 49.4 23 54 200 387 15.0 25.9 9.8e-1 -
7VX690T ResNet-34 i8,i32 7.33G 21.3 23 54 200 182 15.0 12.2 4.0e-2 -
7VX690T VGG16 i8,i32 30.6G 14.7 23 54 200 344 15.0 23.0 8.9e-2 -
(Wei et al., 2019) 7K325T LeNet-5 f32 f32 - 6.6 34 33 100 - - - 2.3e-3 -
7K325T LeNet-5 i8 i8 - 1.7 14 16 100 - - - 2.3e-3 -
(Yan et al., 2022) 7A200T Impr. VGG16 i8 40.9G 14.8 12 29 200 23.1 3.4 6.8 1.8e+0 -
7A200T Impr. YOLOv2 i8 379 G 49.4 12 29 200 22.2 3.4 6.5 1.7e+1 -
(Zhang et al., 2021) 7Z035 Impr. YOLOv2 i8,f32 379 G - 21 74 200 111.5 6.0 18.7 3.4e+0 -
ML S (Fraczek et al., 2018) 7Z020 Decision Tree i8 - - 10 86 - - 3.5 - 3.0e-1* -
(Martins et al., 2024) 7Z020 SVM f32* - - 38 - 200 - - - 1.1e-4* -
@@
(Ratnakumar and Nanda, 2021) 6VLX240T RDBC - - - - - 230 - 0.1 - 2.9e-2 -
@@
HLS CNN S (Kim et al., 2024) 7Z020 TriCloudNet + U-Net - - 3.6 49 60 - - 0.2 - 1.0e+0 1
(Wang and Qiu, 2019) 7Z020 CNN b - - 11 100 - 69.8 2.4 29.3 - 7
(Yang et al., 2022) 7VX690T CNN2@0.7 i? 450 M 0.1 64 21 250 298 5.8 51.4 7.0e-4 652
7VX690T CNN4@0.7 i? 560 M 0.1 69 21 250 366 6.6 55.5 7.0e-4 636
7VX690T CNN6@1.6 i? 600 M 0.1 76 18 250 235 5.1 46.1 1.6e-3 384
F (Wu et al., 2021) ZU9EG Impr. YOLOv3 i8 - 8.5 - - - - - - 7.0e-3 -

### v2

@@ A Large Catalog of Computing Platforms
Through the motivations of the surveyed studies, and in line with the conclusions of Wang et al. (2020) and Lentaris et al. (2018), a generic guideline for hardware platform selection, considering development effort and performance targets, is reported in Table 3. GPUs, TPUs, and VPUs offer quicker paths to achieving initial performance, especially for computationally complex DNNs (Bayer et al., 2024). FPGAs, however, promise superior performance and efficiency if more development effort and hardware expertise are invested in custom accelerator design and optimization. For applications demanding ultimate performance and willing to bear the highest development costs and longest design cycles, custom ASICs represent the pinnacle, albeit with reduced updatability (Amara et al., 2006). In addition, both ASICs and FPGAs can directly connect their compute units to input and output ports, enabling lower latencies compared to GPUs and CPUs. In the following sections, we examine the technical aspects of the surveyed studies, exploring the methodologies and techniques employed to implement and optimize ML models on FPGAs for RS applications.
AMD FPGAs overwhelmingly dominate the surveyed hardware, comprising nearly all devices ($96\%$). The preferred FPGAs are SoC boards, particularly the Zynq family ($57\%$). Most boards are equipped to run DNNs ($60\%$ with >500>500 DSPs), and almost all devices have accelerated at least one DL model in an experiment. Reconfigurability, power efficiency, and radiation concerns are the primary motivations for choosing FPGAs over other DNN-friendly COTS accelerators. 2.1 TAKEAWAYS
@@ 4. From AI Models to FPGA Deployment: Design Strategies and Challenges
Before deploying a model on an FPGA, it typically requires adaptation and compression. These optimization steps are closely linked to hardware design decisions—discussed later in  2.1 and  2.1. In  2.1, we introduce key network compression techniques and discuss their impact and requirements. Since traditional ML methods are already lightweight, we primarily discuss optimizations for computationally intensive models, like NNs.
@@
Most NNs include a feature extractor, or backbone, responsible for refining and compressing the input’s essential information. Extracting rich, high-quality features often relies on large and deep backbones, which come with high computational costs. To lighten these costs, designers can exploit lightweight backbones. For instance, Yang et al. (2023) replace YOLOv3’s DarkNet53 backbone with GhostNet (Han et al., 2020), reducing the number of learnable parameters by a factor of $3$. Similarly, Li et al. (2023a) swap YOLOv4’s CSPDarknet53 backbone for MobileNeXt (Zhou et al., 2020), lowering the parameter count from 60.0860.08M to 36.4436.44M, with only a $1.65\%$ mAP drop—compensated via other customizations. As discussed in  2.1, the existence of many architectures make selection a challenging task. Sabogal and George (2021) compare four DL segmentation models, showing that ENet (Paszke et al., 2016) outperforms U-Net (Ronneberger et al., 2015) while being $23.81\times$ more computationally efficient and requiring $20.56\times$ less memory. Similarly, Yang et al. (2022) compare three popular lightweight backbones: MobileNetv1 (Howard et al., 2017), MobileNetv2 (Sandler et al., 2018), and SqueezeNet (Iandola et al., 2016), concluding that MobileNetv1 achieves the best trade-off for their SAR ship detection application. MobileNetv1 also introduces a width multiplier for layer-wise scaling, leveraged by Suh et al. (2021) to explore different deployment scenarios in terms of accuracy and FPGA resource use. Beyond swapping architectures, Kim et al. (2024) propose a multi-phase cloud coverage detection pipeline where images first undergo a uniformity check and classification via a small CNN (TriCloudnet) before full segmentation with U-Net. Such workflow avoids the computation of cloud-clear or overcast images, reducing processing time and power consumption by $48.7\%$. When sacrificing some feature representation capabilities is manageable, lightweight and multi-phase approaches highlight that questioning a model’s depth and width can result in significant computational savings.
@@
Given the ubiquity of convolution operations in Computer Vision, lightweight backbones commonly focus on efficiently computing them. For instance, MobileNets architectures (Howard et al., 2017; Sandler et al., 2018; Howard et al., 2019) replace standard convolution layers by depth-wise separable convolutions. This form of factorized convolutions separate the standard convolution operation in two stages: filtering, with single 2D filters for each input channel, and combination, by linearly combining the output of the filtering with a 1x1 convolution. When considering an input tensor $h_{i}\cdot w_{i}\cdot d_{i}$, a kernel $k\cdot k\cdot d_{i}\cdot d_{j}$, and an output tensor $h_{i}\cdot w_{i}\cdot d_{j}$, a standard convolution costs $h_{i}\cdot w_{i}\cdot d_{i}\cdot d_{j}\cdot k\cdot k$, while a depth-wise separable convolution costs $h_{i}\cdot w_{i}\cdot d_{i}(k^{2}\cdot d_{j})$, a computational reduction of a factor $\approx k^{2}$. For example, Howard et al. (2017) use depth-wise separable convolution to reduce computation by $9\times$ for a loss of $1\%$ on ImageNet. As an alternative path to convolution approximation, All Adder Networks (Chen et al., 2020a) replace the multiplications in convolution kernels by additions. While it does not reduce computational complexity, additions are generally less expensive than multiplications. In particular, the All Adder NN of Zhang et al. (2023) achieves $3\text{\,}$ TOP/s, the highest computational throughput of the survey.
@@
Pruning reduces computational complexity and memory costs by removing non-essential weights or entire channels from an NN. While the usual overparameterization of NNs aids training, it results in many redundant weights (Sze et al., 2017). After training, pruning nullifies the least significant weights (in contribution to the output) to reduce this redundancy, but often requires fine-tuning to recover lost accuracy. The resulting trade-off between parameter reduction and task accuracy can be adjusted with a pruning factor, which determines the percentage of weights to be nullified. To further enhance pruning efficiency, sparsity is often enforced during training via techniques like L1-regularization (Zhang et al., 2022; Zhao et al., 2023). For example, Zhang et al. (2022) train their GNN with lasso regression—adding an L1-penalty term to the loss—to encourage weight shrinkage and facilitate pruning. Channel pruning removes entire feature channels, enabling more efficient hardware execution. Yang et al. (2023) prune Ghost-YOLO channels, reducing the number of parameters from $23.52\text{\,}\mathrm{M}$ to $6.71\text{\,}\mathrm{M}$ at a cost of $2.93\text{\,}\%$ mAP [^13] High accuracy costs are usually mitigated with other optimizations, in this case via Knowledge Distillation (KD) (Hinton et al., 2015), discussed further in Section 6.. Similarly, Nguyen et al. (2024) prune $60\text{\,}\%$ of filter weights, incurring only a $0.85\text{\,}\%$ mAP loss. Beyond weights and channels, certain families of models also have the opportunity to prune input data. For example, Zhang et al. (2022) reduce the computational complexity of their GNN by $92.8\text{\,}\%$ by pruning input graph vertices, causing only a $0.17\text{\,}\%$ OA drop. To leverage the efficiency improvements reached through pruning, Yang et al. (2022) perform hardware-guided progressive pruning. The method integrates an initial coarse pruning, a resource-per-layer cost estimation, and a final resource-aware fine pruning. Pruning is frequently paired with quantization for further efficiency gains (Han et al., 2016; Zhao et al., 2023).
@@
One of FPGA’s key computational advantages over alternative hardware platforms, like GPUs, lies in custom datatype operations. Quantization reduces numerical precision by converting full-precision 32-bit floating-point (fp32) parameters into lower-precision formats such as 16-bit fixed-point (i16), 8-bit integer (i8), or even binary (b). Low-bandwidth datatypes enable highly efficient arithmetic operations, particularly on FPGA logic, where binary multiplications can be performed using simple XOR operations. While quantization reduces computational cost and memory footprint, it inevitably perturbs model parameters, [^14] Several factors influence the effectiveness of a quantization scheme, such as its uniformity, symmetry, and granularity. For a more detailed discussion on quantization techniques, we recommend the comprehensive survey of Gholami et al. (2021). shifting them from their convergence point and causing accuracy degradation compared to full-precision implementations (Gholami et al., 2021). For instance, Sabogal and George (2021) report an average accuracy drop of $1\%$ across four segmentation models. Similarly, in object detection, Zhao et al. (2023) observe a $0.42\%$ drop in mAP, while Ni et al. (2023) report a minimal reduction in classification accuracy of $0.03\%$ for VGG16 and $0.06\%$ for ResNet-34. Although the quantization-induced loss rarely exceeds $1\%$, it can be unacceptable in scenarios requiring high model reliability. To mitigate accuracy degradation, designers may opt to re-train the network, effectively choosing one of two standard approaches: Post-Training Quantization (PTQ) or Quantization-Aware Training (QAT). PTQ, sometimes called direct quantization, is straightforward to implement as it applies quantization directly after training. However, PTQ’s simplicity can cause significant accuracy degradation, especially in models with a wide parameter distribution (Jacob et al., 2018). For instance, Wei et al. (2019) observe over $50\%$ accuracy loss across all tested bit widths when applying PTQ to a LeNet-5 for MSTAR classification. In contrast, QAT simulates low-precision arithmetic already during training, resulting in better robustness to quantization noise (Pitonak et al., 2022), albeit at a higher computational cost (Yu et al., 2024). Beyond PTQ and QAT, mixed-precision quantization assigns different bit widths to different sections of the model. This method preserves accuracy by keeping sensitive layers at higher precision while applying aggressive quantization to less critical layers. For example, Yang et al. (2022) identify inefficient layers by estimating layer-wise latency and resource consumption. Rapuano et al. (2021) take this idea further by manually assigning custom bit widths to each input, filter, and output, achieving a $48\%$ memory footprint reduction for CloudScout. Alternatively, some approaches like Zhang et al. (2020) mitigate accuracy loss by selectively retaining fp32 precision for activations while quantizing only the weights. Similarly, Wei et al. (2019) introduce a symmetric QAT method that quantizes weights but preserves normalization and activation in fp32 format. The resulting computation overhead is motivated by the need to preserve the model’s accuracy. Quantization is indispensable to achieve FPGA designs that are competitive with higher-frequency computing platforms. While PTQ is often sufficient to achieve 8-bit quantization with minimal accuracy degradation (Nagel et al., 2021), QAT is preferable when maintaining accuracy is critical. Mixed-precision quantization further optimizes efficiency by balancing accuracy and computational cost.
@@
All the methods discussed above effectively optimize NNs, but their combined use often yields the best results. Notably, combining lightweight backbones and fixed-point quantization is a recurring and effective strategy in the surveyed studies (Yang et al., 2023, 2022; Yu et al., 2024; Li et al., 2023a, 2019b; Neris et al., 2022; Nguyen et al., 2024; Zhao et al., 2023; Kim et al., 2024). Additionally, lightweight backbone replacement or the direct use of efficient architectures, such as ENet (Paszke et al., 2016) or “tiny” variants like YOLOv4-tiny (Yu et al., 2024; Nguyen et al., 2024), appear in $\sim\!20\%$ of the surveyed experiments, offering an effective trade-off between accuracy and complexity. When the accuracy drop must be compensated, designers typically enhance the baseline architecture or training setup. For example, Nguyen et al. (2024) add a third YOLOv4 head to detect targets at multiple spatial scales and optimize training with data augmentation to recover accuracy losses. Several surveyed studies (Zhao et al., 2023; Nguyen et al., 2024; Han et al., 2016) combine backbone replacement, pruning, quantization, and a recovery method in an incremental optimization strategy. Such gradual development approaches allow designers to evaluate the contribution of each step in an ablative manner, progressively refining both accuracy and resource efficiency before final FPGA deployment. Ultimately, the effectiveness of these optimization techniques depends not only on compression gains but also on their alignment with FPGA accelerator design, discussed in the next section.
To reduce computational load, 24% of DNNs leverage lightweight backbones or efficient architectures, primarily MobileNets (46%). Out of all DL experiments, 89% use quantization, mostly to int88 format, and 26% leverage pruning to compress their model. Fine-granularity mixed-precision quantization presents significant opportunities for further network compression. 2.1 TAKEAWAYS
@@
Implementing NNs on CPUs and GPUs typically involves using stable and mature software frameworks and drivers. In contrast, automatic deployment frameworks for FPGAs have not seen the same level of adoption—most implementations depend on manual frameworks and varying design patterns. To support future work in selecting suitable implementation frameworks and design patterns, we provide a dedicated second taxonomy in Table 4. It presents key metrics relevant to the respective FPGA implementation, as reported in the original study, organized into three column groups. First, Implementation Choices includes the implementation framework ’Impl.’, the model family ’Fam.’, the design pattern ’P’ (S for Specific, F for Flexible), the ’FPGA’ product, and the ’Model Name’. Second, Design Metrics covers the precision post quantization ’Prec.’, the model complexity ’C[OP], and the memory footprint [MB] (: model parameters are located on-chip, and not loaded from external memory). Additionally, we report the DSP ’D[%] and BRAM ’B[%] resource utilization. Finally, Performance Metrics attempts to compare different designs with each other. In particular, we report the operating frequency in [MHz], computational throughput ’T[GOP/s], power consumption ’P[W], computational efficiency ’T/P’, Latency ’L[s], and temporal throughput in [FPS].
@@
Manual HDL CNN S (Ieracitano et al., 2024) 7Z045 HO-ShipNet i16 - - 44 44 - - 1.90 - 1.01 ms -
(Rapuano et al., 2021) ZU7EV CloudScout i? - 13.30 67 46 115 - 3.40 - 141.7 ms -
(Vitolo et al., 2024) 7A35T CNN f32* - - - - 36 1.1 0.116 9.8 - -
(Yang et al., 2023) 7Z020 Ghost-YOLOS i16 8.41G 12.60 58 60 150 29.5 2.98 9.9 320 ms -
(Zhang et al., 2020) 7VX690T Q-IORN i8 - 121.51 21 27 200 209.6 6.32 33.2 147.62 ms 6.77
(Zhang et al., 2023) 7VX690T A2NN i8,f32 14.9G - 9 86 200 3047 8.27 368.4 4.92 ms 203.25
F (He et al., 2023) ZU15EG 2D CNN i8 684 K 1.20 5 71 100 7.1 8.40 0.8 0.097 ms* -
ZU15EG 3D CNN i8 4.23M 0.12 9 76 100 3.8 8.40 0.5 1.11 ms* -
ZU15EG HybridSN i8 101 M 20.50 37 80 100 13.2 8.40 1.6 7.71 ms* -
(Li et al., 2019b) 7Z100 CBFF-SSD i16 - - 57 60 200 452.8 19.52 23.2 42.59 ms -
(Li et al., 2020b) ZU3EG AP2D-Net i(mix) - - 80 75 300 130.2 5.59 23.3 32.8 ms 30.53
(Ni et al., 2023) 7VX690T Improved YOLOv2 i8,i32 379 G 49.40 23 54 200 387 14.97 25.9 981.4 ms -
7VX690T ResNet-34 i8,i32 7.33G 21.29 23 54 200 182 14.97 12.2 40.2 ms -
7VX690T VGG16 i8,i32 30.6G 14.70 23 54 200 344 14.97 23.0 89.1 ms -
(Wei et al., 2019) 7K325T LeNet-5 f32 f32 - 6.64 34 33 100 - - - 2.29 ms -
7K325T LeNet-5 i8 i8 - 1.66 14 16 100 - - - 2.29 ms -
(Yan et al., 2022) 7A200T Improved VGG16 i8 40.9G 14.80 12 29 200 23.1 3.41 6.8 1780 ms -
7A200T Improved YOLOv2 i8 379 G 49.40 12 29 200 22.2 3.41 6.5 17120 ms -
(Zhang et al., 2021) 7Z035 Improved YOLOv2 i8,f32 379 G - 21 74 200 111.5 5.96 18.7 3400 ms -
ML S (Fraczek et al., 2018) 7Z020 Decision Tree i8 - - 10 86 - - 3.50 - 300 ms* -
(Martins et al., 2024) 7Z020 SVM f32* - - 38 - 200 - - - 0.11 ms* -
@@
(Ratnakumar and Nanda, 2021) 6VLX240T RDBC - - - - - 230 - 0.12 - 28.9 ms -
@@
HLS CNN S (Kim et al., 2024) 7Z020 TriCloudNet + U-Net - - 3.61 49 60 - - 0.196 - 1028 ms 0.97
(Wang and Qiu, 2019) 7Z020 CNN b - - 11 100 - 69.8 2.38 29.3 - 7
(Yang et al., 2022) 7VX690T CNN2@0.7 i? 450 M 0.09 64 21 250 298 5.80 51.4 0.7 ms 652
7VX690T CNN4@0.7 i? 560 M 0.11 69 21 250 366 6.60 55.5 0.7 ms 636
7VX690T CNN6@1.6 i? 600 M 0.09 76 18 250 235 5.10 46.1 1.6 ms 384
F (Wu et al., 2021) ZU9EG Improved YOLOv3 i8 - 8.50 - - - - - - 7 ms -